# Nanocavities for Molecular Optomechanics: their fundamental description and applications

### v1

@@ Conclusion
Looking ahead, we can see numerous promising and challenging research directions. Hybrid plasmonic-dielectric approaches may allow to reach overcoupled cavities as well as deeply sideband-resolved McOM. New ideas are needed to prevent metal surface restructuring under high laser power, and the failure mode under pulsed excitation should be clarified. Molecules with long-lived vibrational modes and large Raman cross-section should be engineered to boost the cooperativity. Entanglement between plasmon and molecular vibrations, and among vibrating molecules, should be demonstrable. Exploiting resonant or near-resonant coupling to electronic molecular transitions would establish tripartite phonon-photon-exciton systems 147, 148, 149, 150, 136, 151, bearing analogy to macroscopic optomechanical systems that include two-level systems, with potential for enhanced coupling strengths, nonlinear effects, and applications in wavelength conversion 152, 153. Finally, the realization of the single-photon optomechanical strong coupling would give rise to a nonlinear response at the quantum level, both for photonic and phononic degrees of freedom, a holly grail in quantum optomechanics 154, 155.
@@ A Mie approach: spherical nanoparticle
The Mie theory gives an exact solution to the problem of the scattering of an electromagnetic (EM) wave by a sphere. In addition to the EM field solutions at any point in space, this theory conveniently provides analytical expressions for integral quantities such as the scattered power $P_{\mathrm{sca}}$ and the absorbed power $P_{\mathrm{abs}}$ (see Fig. 7). For a sufficiently small silver nanosphere, all parameters of the plasmonic dipole mode can be accurately described using scattering expressions. Using a Mie solver package 1, the spectra $P_{\mathrm{sca}}(\lambda)$, $P_{\mathrm{abs}}(\lambda)$ and $P_{\mathrm{ext}}(\lambda)=P_{\mathrm{sca}}(\lambda)+P_{\mathrm{abs}}(\lambda)$ are obtained by scanning the wavelength of the plane-wave excitation (PWE) at constant incident intensity ($S_{\mathrm{inc}}=P_{\mathrm{inc}}/A_{\mathrm{focus}}$, with $A_{\mathrm{focus}}$ the effective focal area of the incoming light away from the nanoparticle). Analytical expression of $A_{\mathrm{focus}}$ for different light waves and its connection with the normalized energy density of the wave, can be found in Ref. 156.
@@
This simple approach can be applied to more elaborated structures via FEM and FTDT simulations as long as the targeted mode is sufficiently well-isolated (spectrally or spatially from the contributions of other modes).
@@
The total Purcell factor $M_{\mathrm{tot}}$ calculated from the PDE calculations is dominated by a continuum of modes 93 even at the dipolar mode’s resonance frequency so that direct mode volume calculation would not be possible. On the contrary, for a PDE tuned in resonance with a sufficiently isolated cavity mode, the radiative part of the Purcell factor $M_{\mathrm{rad}}$ is, to a very large extent, due to this mode only. With the help of the efficiencies calculated previously, we can approximate the total Purcell factor due to a single dipolar bright mode as: $M^{\mathrm{(dip)}}_{\mathrm{tot}}=M_{\mathrm{rad}}(1+\eta_{\mathrm{abs}}/\eta_{\mathrm{rad}})=M_{\mathrm{rad}}(1+\kappa_{\mathrm{0}}/\kappa_{\mathrm{rad}})$. This factor can be used directly to approximate the mode volume following the approach of Ref. 7 or be linked to the approach exposed above via the following relation between Purcell and field enhancement factors: $K^{2}=\eta_{\mathrm{rad}}M^{\mathrm{(dip)}}_{\mathrm{tot}}$. It has to be noted that both mode volume estimates given here are in very good agreements with the values found by F. Koenderink in Ref. 157 for small spherical Ag NP and with our own evaluations via FEM calculations of the mode volume following a quasi-normal mode (QNM) treatment 83. A similar level of agreement is found for the other cavity parameters calculated for the single NP case.
@@
A rigorous modal treatment, based on the quasi-normal mode formalism, is possible for many of the parameters introduced above 83. The direct evaluation of $\eta_{\mathrm{rad}}$, named modal brightness $\mathcal{B}$ in the QNM framework, remains however a tedious task. Indirect calculations of $\eta_{\mathrm{rad}}$ for NPoM cavities have been introduced in Ref. 87. Performing PDE calculations followed by near-to-far field transformation (NFFT) 158, the energy dissipated through radiative and guided-mode (SPP) channels can be separated. This way, $\eta_{\mathrm{ant}}$ and $\beta\eta_{\mathrm{spp}}$ are found. Additional QNM calculations enable to evaluate $\beta$ so that the multimode absorption channel $\beta\eta_{\mathrm{abs}}=\beta-\eta_{\mathrm{ant}}-\beta\eta_{\mathrm{spp}}$ as well as all single-mode counterparts $\eta_{\mathrm{rad}},\eta_{\mathrm{spp}},\eta_{\mathrm{abs}}$} can be quantified.
In this Perspective, and in analogy with the single NP treatment discussed above, we perform PWE calculations on resonance with the plasmonic mode of interest followed by NFFT to evaluate the single-mode loss channels $\eta_{\mathrm{rad}}$ and $\eta_{\mathrm{spp}}$. Similarly to the treatment of Ref. 157, we evaluate the divergence of the Ohmic losses with the size of the integration domain 159 to separate the SPP contribution to absorption from the single-mode localized absorption ($\eta_{\mathrm{abs}}$).

### v2

@@ Conclusion
Looking ahead, we can see numerous promising and challenging research directions. Hybrid plasmonic-dielectric approaches may allow to reach overcoupled cavities as well as deeply sideband-resolved McOM. New ideas are needed to prevent metal surface restructuring under high laser power, and the failure mode under pulsed excitation should be clarified. Molecules with long-lived vibrational modes and large Raman cross-section should be engineered to boost the cooperativity. Entanglement between plasmon and molecular vibrations, and among vibrating molecules, should be demonstrable. Exploiting resonant or near-resonant coupling to electronic molecular transitions would establish tripartite phonon-photon-exciton systems 148, 149, 150, 151, 137, 152, bearing analogy to macroscopic optomechanical systems that include two-level systems, with potential for enhanced coupling strengths, nonlinear effects, and applications in wavelength conversion 153, 154. Finally, the realization of the single-photon optomechanical strong coupling would give rise to a nonlinear response at the quantum level, both for photonic and phononic degrees of freedom, a holly grail in quantum optomechanics 155, 156.
@@ A Mie approach: spherical nanoparticle
The Mie theory gives an exact solution to the problem of the scattering of an electromagnetic (EM) wave by a sphere. In addition to the EM field solutions at any point in space, this theory conveniently provides analytical expressions for integral quantities such as the scattered power $P_{\mathrm{sca}}$ and the absorbed power $P_{\mathrm{abs}}$ (see Fig. 7). For a sufficiently small silver nanosphere, all parameters of the plasmonic dipole mode can be accurately described using scattering expressions. Using a Mie solver package 1, the spectra $P_{\mathrm{sca}}(\lambda)$, $P_{\mathrm{abs}}(\lambda)$ and $P_{\mathrm{ext}}(\lambda)=P_{\mathrm{sca}}(\lambda)+P_{\mathrm{abs}}(\lambda)$ are obtained by scanning the wavelength of the plane-wave excitation (PWE) at constant incident intensity ($S_{\mathrm{inc}}=P_{\mathrm{inc}}/A_{\mathrm{focus}}$, with $A_{\mathrm{focus}}$ the effective focal area of the incoming light away from the nanoparticle). Analytical expression of $A_{\mathrm{focus}}$ for different light waves and its connection with the normalized energy density of the wave, can be found in Ref. 157.
@@
This simple approach can be applied to more elaborated structures via FEM and FDTD simulations as long as the targeted mode is sufficiently well-isolated (spectrally or spatially from the contributions of other modes).
@@
The total Purcell factor $M_{\mathrm{tot}}$ calculated from the PDE calculations is dominated by a continuum of modes 93 even at the dipolar mode’s resonance frequency so that direct mode volume calculation would not be possible. On the contrary, for a PDE tuned in resonance with a sufficiently isolated cavity mode, the radiative part of the Purcell factor $M_{\mathrm{rad}}$ is, to a very large extent, due to this mode only. With the help of the efficiencies calculated previously, we can approximate the total Purcell factor due to a single dipolar bright mode as: $M^{\mathrm{(dip)}}_{\mathrm{tot}}=M_{\mathrm{rad}}(1+\eta_{\mathrm{abs}}/\eta_{\mathrm{rad}})=M_{\mathrm{rad}}(1+\kappa_{\mathrm{0}}/\kappa_{\mathrm{rad}})$. This factor can be used directly to approximate the mode volume following the approach of Ref. 7 or be linked to the approach exposed above via the following relation between Purcell and field enhancement factors: $K^{2}=\eta_{\mathrm{rad}}M^{\mathrm{(dip)}}_{\mathrm{tot}}$. It has to be noted that both mode volume estimates given here are in very good agreements with the values found by F. Koenderink in Ref. 158 for small spherical Ag NP and with our own evaluations via FEM calculations of the mode volume following a quasi-normal mode (QNM) treatment 83. A similar level of agreement is found for the other cavity parameters calculated for the single NP case.
@@
A rigorous modal treatment, based on the quasi-normal mode formalism, is possible for many of the parameters introduced above 83. The direct evaluation of $\eta_{\mathrm{rad}}$, named modal brightness $\mathcal{B}$ in the QNM framework, remains however a tedious task. Indirect calculations of $\eta_{\mathrm{rad}}$ for NPoM cavities have been introduced in Ref. 87. Performing PDE calculations followed by near-to-far field transformation (NFFT) 159, the energy dissipated through radiative and guided-mode (SPP) channels can be separated. This way, $\eta_{\mathrm{ant}}$ and $\beta\eta_{\mathrm{spp}}$ are found. Additional QNM calculations enable to evaluate $\beta$ so that the multimode absorption channel $\beta\eta_{\mathrm{abs}}=\beta-\eta_{\mathrm{ant}}-\beta\eta_{\mathrm{spp}}$ as well as all single-mode counterparts $\eta_{\mathrm{rad}},\eta_{\mathrm{spp}},\eta_{\mathrm{abs}}$} can be quantified.
In this Perspective, and in analogy with the single NP treatment discussed above, we perform PWE calculations on resonance with the plasmonic mode of interest followed by NFFT to evaluate the single-mode loss channels $\eta_{\mathrm{rad}}$ and $\eta_{\mathrm{spp}}$. Similarly to the treatment of Ref. 158, we evaluate the divergence of the Ohmic losses with the size of the integration domain 160 to separate the SPP contribution to absorption from the single-mode localized absorption ($\eta_{\mathrm{abs}}$).